50.2 kilograms expressed in micrograms is (choose answer with correct significant digits):

a) 5.02 x 107  micrograms

b) 5.02 x 104  micrograms

c) 5.02 x 1010  micrograms

d) 5.02 x 10-2 micrograms
Answer #1

Answer: c) 5.02 ×10^10 micrograms(μg)

1 kg = 10^9 μg

50.2 Kg = 50.2 ×10^9 μg = 5.02 ×10^10 μg

**feel free to comment in any kind of doubt and give thumbs up.